# InterMT: Multi-Turn Interleaved Preference Alignment with Human Feedback

### v1

@@ G.1 Review of Human Annotation Dimensions
G.2 Judge Settings and Metrics
The dataset includes multi-turn multimodal interleaved communication histories and human-annotated ground truth. Evaluated models must assess the conversation at both the turn and conversation levels across nine dimensions, following a set of guidelines. Scoring Evaluation requires the model to assign scores on a 0-3 scale, with evaluation based on agreement and Pearson similarity [109, 52, 44]. Pair Comparison directly compares two individual turns or entire conversations, without considering ties, and is evaluated for accuracy against human judgments. Crucial Step Recognition addresses a key challenge in multi-turn conversations: accurately identifying the user’s intent and determining whether it has been fulfilled, evaluated by the score provided by judge according to the human-annotated reference answers.
@@ G.3 MLLM as a Judge
Inspired by [44, 52], we leverage genuine human-annotated data collected in InterMT to construct InterMT-Bench, a benchmark designed to evaluate the alignment between models and human values in multi-turn multimodal interaction scenarios. Our evaluation focuses on three key aspects: Score Evaluation, Pair Comparison, and Crucial Step Recognition. The system and user prompts used for Score Evaluation and Pair Comparison are illustrated in Figure 15, 17 and Figure 19,21, respectively.
@@
System Prompt: You are a scoring model for evaluating the overall quality in multi-turn visual dialogues. You will receive a conversation history, please read it carefully. Next, I will provide you with an evaluation list and corresponding scoring criteria. Please score the conversation history based on the scoring criteria. Your output needs to be: $[\text{Evaluation Criterion}_{1},\ \boxed{\text{score1}}],\ [\text{Evaluation Criterion}_{2},\ \boxed{\text{score2}}],\ \ldots$ Example: Evaluation list: [context_awareness, helpfulness, crucial_step_recognition, global_image_text_consistency, style_coherence] Output: [context_awarenessscore1], [helpfulnessscore2], [crucial_step_recognitionscore3], [global_image_text_consistencyscore4], [style_coherencescore5] <Annotation Documents> User Prompt: Now, please evaluate the conversation history based on the scoring criteria. And output the result in the format of: $[\text{Evaluation Criterion}_{1},\ \boxed{\text{score1}}],\ [\text{Evaluation Criterion}_{2},\ \boxed{\text{score2}}],\ \ldots$
@@
System Prompt: You are a scoring model for evaluating the quality of a single turn in multi-turn visual dialogues. You will receive a conversation history, please read it carefully. Next, I will provide you with an evaluation list and corresponding scoring criteria. Please score the conversation history based on the scoring criteria. Your output must follow this exact format: Evaluation list: [local_image_text_consistencyvisual_perceptual_qualitytext_qualitycontext_coherence] Output: [[local_image_text_consistencyscore]], [[visual_perceptual_qualityscore]], [[text_qualityscore]], [[context_coherencescore]] Where score is your numerical rating (0–3). <Annotation Documents> User Prompt: Now, please evaluate this turn based on the scoring criteria. Your score should be between 0 and 3. And output the result in the format: $[\![\text{Evaluation Criterion}_{1},\ \boxed{\text{score1}}],\ [\text{Evaluation Criterion}_{2},\ \boxed{\text{score2}}],\ \ldots\!]$
@@
System Prompt: You are a scoring model for evaluating the quality of a single turn in multi-turn visual dialogues. You will receive a conversation history, please read it carefully. Next, I will provide you with an evaluation list and corresponding scoring criteria. Please score the conversation history based on the scoring criteria. Your output must follow this exact format: Evaluation list: [local_image_text_consistencyvisual_perceptual_qualitytext_qualitycontext_coherence] Output: [[local_image_text_consistencyreasonscore]], [[visual_perceptual_qualityreasonscore]], [[text_qualityreasonscore]], [[context_coherencereasonscore]] Where score is your numerical rating (0–3) and reason is your brief justification. <Annotation Documents> User Prompt: Now, please evaluate this turn based on the scoring criteria. Your score should be between 0 and 3. And output the result in the format: $[\![\text{Evaluation Criterion}_{1},\ \text{Reason},\ \boxed{\text{score1}}],\ [\text{Evaluation Criterion}_{2},\ \text{Reason},\ \boxed{\text{score2}}],\ \ldots\!]$
@@
System Prompt: You are a judge model for evaluating the overall quality in multi-turn visual dialogues. You will receive a conversation history, please read it carefully. Next, I will provide you with an evaluation list and corresponding scoring criteria. Please compare the two responses (ResponseA and ResponseB) and give your final preference. Your output needs to follow the format: $[\text{Evaluation Criterion}_{1},\ \boxed{\text{ResponseA}}],\ [\text{Evaluation Criterion}_{2},\ \boxed{\text{ResponseB}}],\ \ldots$ <Annotation Documents> User Prompt: Now, please evaluate the conversation history based on the scoring criteria. And output the result in the format: $[\text{Evaluation Criterion}_{1},\ \boxed{\text{ResponseA}}],\ [\text{Evaluation Criterion}_{2},\ \boxed{\text{ResponseB}}],\ \ldots,$
@@
System Prompt: You are a scoring model for evaluating the overall quality in multi-turn visual dialogues. You will receive a conversation history, please read it carefully. Next, I will provide you with an evaluation list and corresponding scoring criteria. Please score the conversation history based on the scoring criteria and provide a reason for your score. Your output needs to follow the format: $[\text{Evaluation Criterion}_{1},\ \text{Reason},\ \boxed{\text{ResponseA}}],\ \ldots,$ <Annotation Documents> User Prompt: Now, please evaluate the conversation history based on the scoring criteria. And output the result in the format: $[\text{Evaluation Criterion}_{1},\ \text{Your Judge Reason},\ \boxed{\text{ResponseA}}],\ \ldots,$
@@
Table 7: Human agreement percentage with different judge models. Each judgment is independently reviewed by different annotators.
@@ G.4 Details of Crucial Step Recognition Evaluation
During evaluation, we collect human-annotated rationales for crucial step recognition as reference answers. Given recent findings that advanced models can achieve human-comparable performance in pairwise response comparison [52, 44], we employ GPT-4o [38] as the judge to perform partial order comparisons between model outputs (Figure 23 presents the system and user prompts for evaluated models) and the reference answers (Figure 24 presents the system and user prompts for judge models). Each judgment is subsequently reviewed by three human experts, and only those achieving a predefined agreement threshold are considered valid. Final scores are computed by averaging across all evaluation points.
@@ G.5 More Results
Reasoning Ability Is Not a Panacea.
We compare weak reasoning (i.e., providing plausible explanations for the evaluation results) with strong reasoning (i.e., using advanced reasoning models like o4-mini) on the scoring evaluation and pair comparison settings. However, the results are suboptimal. The models’ reasoning processes are primarily based on a step-by-step comparison against predefined guidelines, rather than actively identifying potential flaws in the responses. This approach, which differs from the more granular feedback humans provide, leads to misalignments with human judgment.
Divide and Conquer is beneficial for Crucial Step Recognition
We observe that models with high scores in Crucial Step Recognition tend to adopt a divide-and-conquer approach, meaning they first assess each turn in a multi-turn, multimodal dialogue for problem-solving and alignment with human intent, and then provide an overall conclusion; in contrast, models with lower scores often give more generalized responses.
@@ H.1 Preliminaries of Preference Modeling
A widely adopted approach for modeling human preferences is to employ a preference predictor grounded in the Bradley–Terry (BT) model [48]. Given a pair of answers $({\bm{y}}_{1},{\bm{y}}_{2})$ generated from an question ${\bm{x}}$, BT model indicates that the human preference distribution $p^{*}$ [15, 17] can be expressed based on the underlying human reward function $r^{*}({\bm{y}},{\bm{x}})$ as
@@
However, when extending to multi-turn settings, new challenges arise—particularly in capturing the dynamics of evolving user preferences across turns. Moreover, traditional outcome-level reward signals often fail to generalize in purely textual domains [49], let alone in complex multimodal settings involving interleaved understanding and generation. InterMT incorporates both local and global human annotations in multi-turn, multimodal interactions, leading us to investigate efficient preference modeling methods under this more realistic and challenging scenario.
@@ H.2 Long Horizon Human Value Preference Modeling
Inspired by [50, 51], we investigate two strategies for modeling long-horizon preferences in multi-turn multimodal scenarios: prefix preference and chain-based preference. Let $\mathcal{D}_{\text{multi-turn}}=\{({\bm{x}}_{1}^{(i)},{\bm{y}}_{1}^{(i)},\cdots,{\bm{x}}_{k_{i}}^{(i)},{\bm{y}}_{k_{i}}^{(i)})\}_{i=1}^{N}$ denote the multi-turn human image-text dataset, where $k_{i}$ denotes the number of turns for each conversation. The prefix-preference approach models preferences at the turn level. Given a prefix of the conversation history, it aims to identify the preferred candidate response for the current turn, effectively capturing fine-grained local preferences. The training objective for the prefix-preference reward model $r_{\phi_{\text{prefix}}}({\bm{y}},{\bm{x}})$ is

### v2

@@ G.1 Review of Human Annotation Dimensions
G.2 Judge Settings and Evaluation Metrics
The dataset includes multi-turn multimodal interleaved communication histories and human-annotated ground truth. Evaluated models must assess the conversation at both the turn and conversation levels across nine dimensions, following a set of guidelines. Scoring Evaluation requires the model to assign scores on a 0-3 scale, with evaluation based on agreement and Pearson similarity [lee1988thirteen, zheng2023judging, chen2024mllm]. Pair Comparison directly compares two individual turns or entire conversations, without considering ties, and is evaluated for accuracy against human judgments. Crucial Step Recognition addresses a key challenge in multi-turn conversations: accurately identifying the user’s intent and determining whether it has been fulfilled, evaluated by the score provided by judge according to the human-annotated reference answers.
@@ G.3 MLLM as a Judge
Inspired by [chen2024mllm, zheng2023judging], we leverage genuine human-annotated data collected in InterMT to construct InterMT-Bench, a benchmark designed to evaluate the alignment between models and human values in multi-turn multimodal interaction scenarios. Our evaluation focuses on three key aspects: Score Evaluation, Pair Comparison, and Crucial Step Recognition. The system and user prompts used for Score Evaluation and Pair Comparison are illustrated in Figure 15, 17 and Figure 19,21, respectively.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Table 9: Supplementary Score Evaluation of InterMT-Bench: Human Agreement. Percentage of agreement between human annotators and different judge models. Each judgment is independently reviewed by multiple annotators.
@@ G.4 Details of Crucial Step Recognition Evaluation
During evaluation, we collect human-annotated rationales for crucial step recognition as reference answers. Given recent findings that advanced models can achieve human-comparable performance in pairwise response comparison [zheng2023judging, chen2024mllm], we employ GPT-4o [openai2024gpt4o] as the judge to perform partial order comparisons between model outputs (Figure 23 presents the system and user prompts for evaluated models) and the reference answers (Figure 24 presents the system and user prompts for judge models). Each judgment is subsequently reviewed by three human experts, and only those achieving a predefined agreement threshold are considered valid. Final scores are computed by averaging across all evaluation points.
@@ G.5 More Results
Pairwise comparison evaluated using spearman correlation
We thank the reviewer HqHt for the careful reading, statistical insights, and constructive feedback. Inspired by the suggestion, we conduct a detailed pairwise comparison of model performance on InterMT-Bench using Spearman correlation. Table 10 summarizes the results.
Table 10: InterMT-Bench pairwise comparison results (metric: spearman). L1–L4 and G1–G5 denote different local and global evaluation subsets, respectively.
Model L1 L2 L3 L4 Avg. G1 G2 G3 G4 G5 Avg.
GPT-4.1 0.1719 0.1365 0.1853 0.1562 0.1625 0.5592 0.5281 0.5446 0.4618 0.3111 0.4809
GPT-4.1 (+reason) 0.2175 0.1725 0.1482 0.1893 0.1819 0.3600 0.4543 0.4626 0.4710 0.0449 0.3586
GPT-4o 0.1192 0.0486 0.0889 0.1203 0.0942 0.3304 0.4550 0.3622 0.3193 0.2440 0.3422
GPT-4o (+reason) 0.0674 0.1014 0.0381 0.0864 0.0733 0.1341 0.2395 0.2472 0.2235 0.0655 0.1820
Gemini-Pro* 0.0787 0.0749 0.1354 0.0126 0.0754 0.5019 0.4662 0.2890 0.1052 0.0682 0.2861
Gemini-Pro* (+reason) 0.0535 0.0247 0.1055 0.0009 0.0461 0.4476 0.4188 0.4189 0.3228 0.1841 0.3585
Claude-thinking* 0.1562 0.2247 0.1345 0.1357 0.1628 0.5978 0.5614 0.5438 0.3720 0.1579 0.4466
Claude-thinking* (+reason) 0.1521 0.2007 0.1916 0.1585 0.1757 0.4779 0.4996 0.5364 0.4838 0.1421 0.4280
o4-mini 0.0291 0.1736 0.1311 0.1621 0.1240 0.5761 0.5310 0.5392 0.4415 0.2326 0.4641
o4-mini (+reason) 0.1234 0.1515 0.0962 0.1291 0.1250 0.4269 0.4725 0.5293 0.4998 0.1778 0.4213
@@ H.1 Preliminaries of Preference Modeling
A widely adopted approach for modeling human preferences is to employ a preference predictor grounded in the Bradley–Terry (BT) model [bradley1952rank]. Given a pair of answers $({\bm{y}}_{1},{\bm{y}}_{2})$ generated from an question $\bm{x}$}, BT model indicates that the human preference distribution $p^{*}$ [bai2022training, rafailov2024direct] can be expressed based on the underlying human reward function $r^{*}({\bm{y}},{\bm{x}})$ as
@@
However, when extending to multi-turn settings, new challenges arise—particularly in capturing the dynamics of evolving user preferences across turns. Moreover, traditional outcome-level reward signals often fail to generalize in purely textual domains [shani2024multi], let alone in complex multimodal settings involving interleaved understanding and generation. InterMT incorporates both local and global human annotations in multi-turn, multimodal interactions, leading us to investigate efficient preference modeling methods under this more realistic and challenging scenario.
@@ H.2 Long Horizon Human Value Preference Modeling
Inspired by [qiu2024reward, liao2025tpo], we investigate two strategies for modeling long-horizon preferences in multi-turn multimodal scenarios: prefix preference and chain-based preference. Let $\mathcal{D}_{\text{multi-turn}}=\{({\bm{x}}_{1}^{(i)},{\bm{y}}_{1}^{(i)},\cdots,{\bm{x}}_{k_{i}}^{(i)},{\bm{y}}_{k_{i}}^{(i)})\}_{i=1}^{N}$ denote the multi-turn human image-text dataset, where $k_{i}$ denotes the number of turns for each conversation. The prefix-preference approach models preferences at the turn level. Given a prefix of the conversation history, it aims to identify the preferred candidate response for the current turn, effectively capturing fine-grained local preferences. The training objective for the prefix-preference reward model $r_{\phi_{\text{prefix}}}({\bm{y}},{\bm{x}})$ is
@@ H.3 Evaluation Details
Appendix I Application
I.1 Application in Post-training Methods
Table 12: InterMT-enhanced alignment improves single-turn and multi-turn multimodal understanding and generation.
Task MMIE (Single-turn) Interactor (Ours, Multi-turn)
Models Situational Analysis Project-based Learning Multi-step Reasoning Weighted Average Local-Preference Global-Preference Average
Closed-source Models
GPT4o+SD-XL 59.2% 68.6% 49.7% 59.8% 77.3% 56.3% 66.8%
GPT4o+SD-3 58.4% 68.7% 48.9% 59.3% 75.5% 54.8% 65.2%
GPT4o+Flux 58.2% 69.2% 48.7% 59.4% 77.7% 54.5% 66.1%
Claude+SD-XL 53.9% 68.6% 48.3% 57.6% 77.2% 61.0% 69.3%
Claude+SD-3 53.5% 68.6% 48.5% 57.5% 72.8% 59.2% 66.0%
Claude+Flux 53.3% 68.8% 48.8% 57.6% 74.5% 60.7% 67.6%
Open-source Models
AA-Chameleon 41.8% 46.3% 59.5% 44.7% 18.2% 16.0% 17.1%
AA-Chameleon+RS 43.4% 45.3% 60.3% 45.4% (+0.7%) 18.7% 17.8% 18.3% (+1.2%)
AA-Chameleon+SFT 46.3% 56.9% 65.6% 56.6% (+11.9%) 22.5% 21.3% 21.9% (+4.8%)
AA-Chameleon+DPO 45.8% 58.5% 59.0% 54.4% (+9.7%) 20.7% 18.5% 19.6% (+2.5%)
Janus-Pro 60.3% 55.3% 43.4% 53.0% 19.5% 18.0% 18.8%
Janus-Pro+RS 61.6% 55.1% 42.2% 53.9% (+0.9%) 25.7% 22.0% 23.8% (+5.0%)
Janus-Pro+SFT 50.2% 57.3% 58.7% 55.4% (+2.4%) 21.0% 18.5% 19.8% (+1.0%)
Janus-Pro+DPO 50.5% 63.0% 50.9% 54.8% (+1.8%) 20.2% 19.5% 19.8% (+1.0%)
Our InterMT dataset is designed to be a versatile resource for enhancing Large Vision-Language Models (LVLMs) through various post-training alignment techniques. To demonstrate its efficacy, we applied InterMT to improve two strong open-source models, AA-Chameleon [ji2024align] and Janus-Pro [chen2025janus], using three distinct alignment methods: Supervised Fine-Tuning (SFT) [ouyang2022training], Rejection Sampling (RS) [touvron2023llama], and Direct Preference Optimization (DPO) [rafailov2024direct]. The results, summarized in Table 12, show that fine-tuning with InterMT improves performance across both single-turn vision-language benchmarks (e.g., MMIE [xia2025mmie]) and complex multi-turn multimodal tasks (e.g., our Interactor benchmark [^3] Interactor is an evaluation set derived from real human data, independent of the training dataset).
We train the models based on the Align-Anything codebase [^4] https://github.com/PKU-Alignment/align-anything, and implement MMIE following the original benchmark protocols. For each sub-metric, evaluation scores are computed as a weighted average according to the number of valid evaluation instances, consistent with the scoring methodology of the original MMIE benchmark.
It is worth noting that, due to the current absence of fully native multimodal input-output foundation models—most existing models support multiple modes such as TI2T, I2T, and T2I, but can activate only a single mode at a time—InterMT may not yet realize its full potential in a single training pass. During fine-tuning, we need to select which mode first to activate to most effectively enhance performance. Nevertheless, as demonstrated by both InterMT-Bench   and the publicly released InterMT-Judge, even in the absence of fully multimodal native models, InterMT enables promising explorations in preference modeling and lays the groundwork for future advances in multi-turn multimodal human preference alignment.